# Rewarding Doubt: A Reinforcement Learning Approach to Confidence Calibration of Large Language Models

### v1

@@ 3 Rewarding Doubt
In the following, we prove that $R(a,\hat{p},j)$ is maximized when $\hat{p}=p^{*}$ and thus optimizes for a perfect calibration:
@@
Let $f(\hat{p})$ be the expected reward for all values of $\hat{p}$ and $p^{*}$:
@@
Since the logarithm of zero is undefined, we set the actual lower limit of the confidence $\hat{p}$ to 0.001 and the upper limit to 0.999. The normalized reward for correct and incorrect answers for each confidence is visualized in Figure 3. Our reward function rewards incorrect responses with a confidence of zero equally to correct responses with a maximum confidence. That means the model could exploit the reward function by always responding with a wrong answer and a confidence of zero. To counteract this possible scenario, we add a small constant value to the reward function if the answer is correct. This does not change the location of the maximum of the reward function and thus does not alter the optimality but discourages the model from systematically responding incorrectly.
4 Experimental Setup
We evaluate our method in both Single-Answer and Multiple-Answer settings. We prompt the model to provide a confidence for each answer as an integer between 0 and 10, which we normalize for the reward calculation. A confidence of 0 is defined as the model being certain that the answer is incorrect, while 10 is defined as the model being certain the answer is correct. We normalize the reward function to the range of $[-1,1]$.
In the Single-Answer setting we train the model on the TriviaQA dataset (Joshi et al., 2017), which contains question-answer-evidence triplets, from which we only use the questions and answers. For generalization experiments, we evaluate our method on CommonsenseQA (Talmor et al., 2019) and MedQA (Jin et al., 2020), which are multiple-choice question datasets in the commonsense and medical domain, respectively. For the Multiple-Answer setting, we train on the QAMPARI dataset (Amouyal et al., 2023), which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought (Xiong et al., 2024), Surrogate Token (Kadavath et al., 2022), Sequence Probability (Huang et al., 2023b), Self-Consistency (Wang et al., 2022) and Trained Probe (Azaria & Mitchell, 2023). We also compare to the non-finetuned base model in a zero-shot manner, using the same prompt as our Rewarding Doubt method and refer to this setup as Verbalize. In the Multiple-Answer setting we compare only to Trained Probe and Sequence probability, as Chain-of-Tought and Self-Consistency have a high runtime in the Multiple-Answer setting and under-performed in the Single-Answer setting.
4.1 Response Generation
To calibrate and reward the model only on the confidences and not the answers we separate generation in two steps. In the Single-Answer-Setting the model is prompted to generate responses in the format "Answer: <answer>, Confidence: <confidence>". In the first generation step an additional stopping criterion is the "Confidence" token, ensuring the model stops before generating the confidence. This response is then used as input to generate the confidence. For multiple answers, the model is prompted to list all its answers together with a confidence score, where each answer-confidence pair has the same format as in the Single-Answer setting. The model first generates all answers and their corresponding confidences. Individual answers and confidences are parsed with regular expressions, and new queries to generate a confidence for each answer are created by appending tokens preceding the previous confidence token.
4.2 Correctness Assessment
4.3 Implementation Details
4.3.1 Training
As base model for our experiments we use Meta-Llama-3-8B-Instruct (Grattafiori et al., 2024). We employ the 4-bit quantized performance-optimized model version by Unsloth AI (Han et al., 2023). Additionally, we apply LoRA fine-tuning (Hu et al., 2022), reducing the amount of trainable parameters from 8,030,265,345 to 41,947,137. For the Single-Answer setting we train the model for two epochs with a learning rate of 1e-5. For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 32,000 steps with a batchsize of eight and a learning rate of 1e-5 with a reward scale of five. On average the model generated approximately 3.4 answers per fact. If the model fails to generate an answer in the specified format, it is penalized with an out-of-format reward of -3. For a correct answer we additionally reward the model with 0.25.
4.3.2 Baselines
For the Sequence Probability, we compute the average probability for each token in the response. In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric (Zhang et al., 2019). For the trained probe (Azaria & Mitchell, 2023), the original study introduced a custom dataset comprising short statements classified as either true or false. The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements. To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers. The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework. For the architecture of the MLP, we employed the same design as Azaria & Mitchell (2023) and train it for four epochs with a learning rate of 1e-4 until convergence. The exact prompts used for each baseline are provided Appendix A.
5 Results and Discussion
This section presents and discusses the key findings of our experiments.
Rewarding Doubt improves calibration significantly across all settings
To evaluate how well our approach improves calibration, we compare it against the zero-shot LLM (Verbalize) and existing methods in both the Single-Answer and Multiple-Answer settings. The comparisons for the Single-Answer setting are provided in Table 1, while those for the Multiple-Answer setting are shown in Table 2. Across all settings, our approach significantly improves the LLM’s confidence calibration compared to zero-shot verbalization and outperforms most baselines, which often completely fail to produce accurate confidence estimates, in both ECE and AUROC. The only exception is the TriviaQA dataset, where the Trained Probe achieves a slightly better ECE score, with both methods achieving near-perfect results. However, our approach achieves a higher AUROC score indicating that its confidences are more effective for distinguishing between correct and incorrect answers compared to the Trained Probe. Since the Trained Probe is the only other trained method, it is not surprising that it outperforms the non-trained baselines. These results further support the claim by Azaria & Mitchell (2023) that the model’s internal state holds information about the truthfulness of statements, which can serve as an indicator of uncertainty. However, the non-fine-tuned model appears to be unable to effectively associate this internal state with uncertainty. Our approach enables the model to make use of this correlation, resulting in a more accurate expression of the probability that a given answer is correct. Notably, our approach consistently outperforms all baselines in the AUROC metrics, even though our reward function only emphasizes calibration accuracy, whereas the Trained Probe is optimized for AUROC. As illustrated in Figure 4, the Rewarding Doubt method leads to much improved calibration curves compared to the base model for both TriviaQA and QAMPARI. The main outlier on QAMPARI at the confidence level of 0.8 is not representative because of a small sample size in this bin.
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.3459 0.5858
Chain-of-Thought 0.3065 0.6379
Surrogate Token 0.3686 0.5923
Sequence Probability 0.3156 0.7804
Self-Consistency 0.1134 0.8213
Trained Probe 0.0189 0.8173
Rewarding Doubt (ours) 0.0226 0.8592
Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting.
(a) Base model
(b) Fine-tuned by Rewarding Doubt
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.5319 0.6047
Sequence probability 0.3230 0.5942
Trained probe 0.2119 0.6481
Rewarding doubt (ours) 0.1061 0.7268
Rewarding Doubt breaks the pattern of overconfidence
Before fine-tuning the base model, we can observe in Figure 5(a) that the model is very overconfident and almost always reports a confidence of eight or higher. This pattern is also observed by Xiong et al. (2024), who hypothesize that this overconfidence comes from the supervised pre-training, where the training data usually expresses high confidence, leading the model to mimic these surface-level patterns instead of developing a real sense of confidence. After fine-tuning the model with our method, as shown in Figure 5(b), it expresses confidence scores distributed over a much wider range, including lower confidence values. This shift indicates that the model is able to overcome the learned pattern of overconfidence and develops an improved sense of confidence.
Rewarding Doubt shows significant generalization capabilities between datasets
To assess the generalization capabilities of our method, we evaluate the model trained on TriviaQA in an out-of-domain setting, using the CommonsenseQA and MedQA datasets. The results are presented in Table 3 and Table 4, respectively. Our approach is significantly better able to generalize than the Trained Probe method, which performs poorly on both datasets and completely fails to generalize to CommonsenseQA. Additionally, our fine-tuned model is significantly better calibrated than the base model on MedQA. However, on CommonsenseQA, the ECE of our approach is slightly worse than that of the base model. This reveals a flaw in the ECE metric. Notably, the ECE does not account for the range of confidence values. For instance, if a model were to consistently predict a confidence of 6 for a dataset with an overall accuracy of 60%, the ECE would be optimal. However, such confidence values would lack practical utility, as they would not enable meaningful discrimination between individual cases. Therefore, a low ECE should not be the sole criterion for evaluating confidence estimation. A model’s ability to discriminate between correct and incorrect answers, as measured by the AUROC score, should also be considered. Consequently, the improvement in the AUROC score compared to the zero-shot verbalization, supports the increased utility of the confidence estimates produced by the fine-tuned model. Overall, these generalization experiments demonstrate that the uncertainty indicators the model learns are not entirely dataset-specific; rather, our method enables the model to develop a more generalized understanding of confidence.
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.2820 0.5425
Trained Probe 0.4819 0.5374
Rewarding doubt (ours) 0.2930 0.6385
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.4480 0.5075
Trained Probe 0.2099 0.5513
Rewarding doubt (ours) 0.1145 0.6649
Rewarding Doubt can generalize between settings
Beyond generalization across datasets, we also assess whether our approach transfers between different experimental settings. When evaluating the model trained in the Single-Answer setting directly on a Multiple-Answer task, it has a considerably better calibration compared to the base model, as seen in Table 5. While it does not achieve the same level of calibration as the model trained directly in the Multiple-Fact setting, this suggests potential benefits in future applications, particulary for improving the calibration of facts in written texts. Training on single facts could already improve the calibration in open-text applications, making it a promising approach for scenarios where structured multi-fact training is impractical.
Training Evaluation ECE ($\downarrow$) AUROC ($\uparrow$)
Base model Single fact 0.5875 0.5787
Single fact Single fact 0.1536 0.7240
Base model Multi fact 0.5319 0.6047
Single fact Multi fact 0.1777 0.6617
Multi fact Multi fact 0.1061 0.7268
Rewarding Doubt is more resource efficient
An advantage of our approach, not visible in the metrics, is the inference speed. While it initially takes more time for training, inference is highly efficient, as only a small constant number of tokens has to be generated to estimate confidence. In constrast, methods like Chain-of-Thought and Self-Consistency need to generate significantly more tokens for their reasoning process, increasing the computational costs and inference time by several orders of magnitude. Our approach introduces no additional overhead, does not depend on external resources, and directly provides confidence information through verbalization. Thus, the time required to express confidence with our method is significantly less than that of previous approaches.
(a) Base model
(b) Fine-tuned by Rewarding Doubt
Rewarding Doubt does not effect answer correctness
Training confidence calibration with our method only targets the uncertainty estimation abilities and does not effect the performance on the task as seen by the accuracy in Table 6. Additionally, our initial hypothesis that the model could learn to generate incorrect answers and label them accordingly with a low confidence score, does not appear to be a problem. As shown in Table 6, there is no significant difference in accuracy, calibration, or AUROC scores with and without an additional reward for correct answers. Given that our method rewards the model on its expressed confidence and not the answer, it is unlikely that the model deliberately generates incorrect responses. Nevertheless, as the inclusion of this additional reward does not negatively impact performance, it can be retained to ensure robust confidence estimation without unintended side effects across different settings.
Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
Base model 0.3459 0.5858 0.6310
Bonus 0.0226 0.8592 0.6309
No bonus 0.0219 0.8562 0.6278
Table 6: Comparison of effect of Rewarding Doubt on accuracy on the TriviaQA dataset.
Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, generalize across tasks, and maintain computational efficiency, making it an effective approach for accurate confidence calibration in language models.
Limitations
6 Conclusion
In this work, we propose Rewarding Doubt, a novel approach that enables LLMs to express confidence in their answers more accurately using natural language. We leverage reinforcement learning with a reward function that incentivizes diverse and well-calibrated confidence expressions. Fine-tuning with our method significantly improves the model’s ability to convey calibrated confidence, effectively reducing the overconfidence commonly observed in LLMs. This not only enhances the trustworthiness in AI-generated responses but also lays the groundwork for more reliable human-AI collaboration, where models can transparently communicate uncertainty, an essential step toward safer and more accountable AI systems.
Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Prompts
Method Prompt
CoT You will get questions. Read the question, analyze step by step, provide your reasoning steps, answer and your confidence in the response. The confidence should be between 0, 1, 2, 3, 4, 5. A value close to 0 means you think there is a high probability that you could be wrong. The closer the value is to 5, the lower you think is the chance that you could be wrong. The output should have the format ’Thoughts: <reasoning steps>, Answer: <just the answer and nothing else>, Confidence: <confidence>’.
[…]
Table 7: The prompts for each method for the open short form questions.
Method Prompt
CoT You will get test questions with possible options. Read the question, analyze step by step, provide your reasoningsteps, answer and your confidence in the response. The confidence should be between 0, 1, 2, 3, 4, 5. A value close to 0 means you think there is a high probability that you could be wrong. The closer the value is to 5, the lower you think is the chance that you could be wrong. The output should have the format ’Thoughts: <reasoning steps>, Answer: <answer_index>, Confidence: <confidence>’ and nothing else.
[…]
Table 8: The prompts for each method for multiple choice questions.
Method Prompt
[…]
Table 9: The prompts for each method for multiple fact questions.

### v3

@@ Appendix B Proof
The proof is analogous to the proof that the logarithmic scoring function is a proper scoring function.
Let $f(\hat{p})=\mathbb{E}[R(a,\hat{p},j)]$ be the expected reward for all values of $\hat{p}$ and $p^{*}$:
@@
As the logarithm of 0 is undefined, we add a small constant $\epsilon$ in the reward function we use for training:
$R(a,\hat{p},j)=\begin{cases}log(\max(\hat{p},\epsilon)),&\text{if $j(a)$ = 1 (correct)}\\ log(\min(1-\hat{p},1-\epsilon)),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
Through this clipping all confidence predictions between 0 and $\epsilon$, and 1 and $1-\epsilon$, respectively, are rewarded equally. This leads to the model not being able to differentiate between confidence estimations within these ranges. We argue this effect is minor for a sufficiently small $\epsilon$ and can be disregarded in practice.
Appendix C Societal Impact
This work introduces a reinforcement learning approach that enables Large Language Models (LLMs) to express calibrated confidence in their factual answers, advancing safe and trustworthy AI deployment. The method improves reliability and uncertainty awareness in LLMs, which is particularly valuable in high-stakes settings such as medicine, law, or customer support, where overconfident errors can have serious consequences. By optimizing a proper scoring rule during training, our method provides a theoretically sound and generalizable mechanism for aligning confidence with factual correctness—supporting human-AI collaboration and informed decision-making. However, expressing numerical confidence may lead users to overly trust AI systems, especially if the model is well-calibrated statistically but still wrong in important individual cases. This risk calls for careful deployment, appropriate user interfaces that contextualize model confidence, and safeguards against overreliance on AI-generated outputs.
Appendix D Asset Details
The TriviaQA dataset [16] is available at https://nlp.cs.washington.edu/triviaqa/ under the Apache License 2.0. The CommonsenseQA dataset [29] (https://www.tau-nlp.sites.tau.ac.il/commonsenseqa) and the MedQA dataset [15] (https://github.com/jind11/MedQA?tab=readme-ov-file) are licensed under the MIT License. The QAMPARI dataset [2] is published at https://samsam3232.github.io/qampari/ and licensed under the CC0 1.0 Universal license. All these licenses allow broad use, distribution and modification. We use the Meta-Llama-3-8B-Instruct [10] language model, provided at https://huggingface.co/unsloth/llama-3-8b-Instruct-bnb-4bit under the Llama 3 Community License Agreement.